# Pushing the frontiers of subexponential FPT time for Feedback Vertex Set

### v1

@@ 3.7. Attempting to build a $t$-uniform partition of $G-M$
(Running time) The set $\mathcal{T}^{\prime}$ can be computed in polynomial time, and finding the subset $X$ and $\mathcal{T}^{\prime}_{X}$ whose existence is proved by 3.13 can again be done in polynomial time by listing the neighborhoods in $M$ of the trees in $\mathcal{T}^{\prime}$. ∎
@@
Let $t$ be an integer. Consider an instance $(G,k,M,\mathcal{H})$ where none of Rules (KR1), (KR2), (KR4) applies, and such that $d_{M}(v)<t$ for every $v\in V(G-M)$. There exists a polynomial time algorithm that given $(G,k,M,\mathcal{H})$ either returns:
  1. (1)
    a subtree $T$ of the forest $G-M$ such that Rule (KR5) applies,
  2. (2)
    a $\widetilde{K_{t,t}}$ such that Rule (BR2) applies, or
  3. (3)
    a $t$-uniform partition $\mathcal{T}$ of the whole forest $G-M$ with $|\mathcal{T}^{+}|<p_{3}(r,s,t)|M|$ ($p_{3}$ being defined in 3.26).
Proof.
We try to construct a partition of $G-M$ fulfilling item 3, and if we fail we prove that we fall into case of item 1 or item 2. We denote $p=p_{3}(r,s,t)$ in this proof.
The construction proceeds from the deepest leaves to the roots of the trees of $G-M$. For this we define $L_{0}$ as the set of roots of the connected components of $G-M$, and more generally $L_{i}$ the set of vertices of depth $i$ in the trees of the forest $G-M$. Let $h$ be the depth of the deepest leaves of the highest trees of $G-M$, and $F_{i}=\bigcup_{j=i}^{h}L_{j}$ the set of vertices with depth at least $i$. Observe that we have $F_{0}=V(G-M)$.
In a first time we construct $\mathcal{T}_{h}$ a $t$-uniform partition of $F_{h}$.
Observe that two vertices of the same depth cannot share an edge, so we have $G[F_{h}]$ an edgeless graph. For constructing the partition we then have no choice but to take each connected component (containing only one vertex) as a tree in $\mathcal{T}$, and more precisely in $\mathcal{T}_{h}^{-}$ as for $v\in F_{h}$ we have $d_{M}(v)<t$ by hypothesis, and so $\mathcal{T}_{h}^{+}=\emptyset$. $\mathcal{T}_{h}$ is then a $t$-uniform partition as there is no edge between two trees of $\mathcal{T}_{h}$, and each tree have degree in $M$ at most $2t$ as wanted (even $t$ as said above).
Now suppose we are given $\mathcal{T}_{i}$ a $t$-uniform partition of $F_{i}$ with $i\geq 1$, such that $|\mathcal{T}_{i}^{+}|<p$. We will now see how to extend this partition to a new partition $\mathcal{T}_{i-1}$ of $F_{i-1}$, or stop the induction and fall into case of item 1 or item 2.
Let $\mathcal{T}_{i-1}=\mathcal{T}_{i}$, $\mathcal{T}_{i-1}^{+}=\mathcal{T}_{i}^{+}$ and $\mathcal{T}_{i-1}^{-}=\mathcal{T}_{i}^{-}$. Let us now modify $\mathcal{T}_{i-1}$ by considering each $v\in L_{i-1}$ as follows.
For any $v\in L_{i-1}$, let $v_{1},\dots,v_{l}$ be its children, which are in the level $L_{i}$. Observe that by definition of a $t$-uniform partition of $F_{i}$, each of these vertices $v_{j}$ are the root of some trees $T(v_{j})$ in $\mathcal{T}_{i}$. Informally, we will try to group some of the $T(v_{j})\in\mathcal{T}_{i}^{-}$ to create a tree in $\mathcal{T}_{i-1}^{+}$. Let $Y^{-}(v)=\{j\mid T(v_{j})\in\mathcal{T}_{i}^{-}\}$. If $d_{M}({v}\cup\bigcup_{j\in Y^{-}(v)}T(v_{j}))<t$, we set $X(v)=Y^{-}(v)$. Otherwise, we define $X(v)$ as an inclusion-wise minimal set $Y^{\prime}(v)\subseteq Y^{-}(v)$ such that $d_{M}\left(\{v\}\cup\bigcup_{j\in Y^{\prime}(v)}T(v_{j})\right)\geq t$. Let
$T(v)=\{v\}\cup\bigcup_{j\in X(v)}T(v_{j}).$
If $|T(v)|>p_{1}(r,\operatorname{\mu}(T))$ we stop and output $T(v)$ (we call this “Case 1”). From 3.20 we get that (KR5) applies as stated in item 1. Let us now assume that $|T(v)|\leq p_{1}(r,\operatorname{\mu}(T))$.
If $d_{M}(T(v))<t$, we remove from $\mathcal{T}_{i-1}^{-}$ all the $T(v_{j})$ for $j\in X(v)$, and add $T(v)$ to $\mathcal{T}_{i-1}^{-}$. We consider $v$ as treated, and proceed with the next vertex $v\in L_{i-1}$.
Otherwise ($d_{M}(T(v))\geq t$), we also remove from $\mathcal{T}_{i-1}^{-}$ all the $T(v_{j})$ for $j\in X(v)$, and we add $T(v)$ to $\mathcal{T}_{i-1}^{+}$. If $|\mathcal{T}_{i-1}^{+}|<p$, then we consider $v$ as treated, and proceed with the next vertex $v\in L_{i-1}$. Otherwise, $|\mathcal{T}_{i-1}^{+}|=p$, then we stop and output $F=\bigcup_{T\in\mathcal{T}_{i-1}}T$ and $\mathcal{T}_{=p}=\mathcal{T}_{i-1}$ (we call this “Case 2”, and we will prove that it matches the requirements of item 2).
This concludes the description of the algorithm that, given $\mathcal{T}_{i}$ a $t$-uniform partition of $F_{i}$ with $i\geq 1$ such that $|\mathcal{T}_{i}^{+}|<p$, either stops in Case 1, stops in Case 2, or compute a set $\mathcal{T}_{i-1}$.
Claim 3.28.
If the algorithm stops in Case 2, then the output matches the requirements of item 2. Otherwise, when the algorithm has processed all the vertices $v\in L_{i-1}$, then $\mathcal{T}_{i-1}$ is a $t$-uniform partition of $F_{i-1}$ with $|\mathcal{T}_{i-1}^{+}|<p$.
Proof.
Let us consider as input $\mathcal{T}_{i}$ a $t$-uniform partition of $F_{i}$ with $i\geq 1$ such that $|\mathcal{T}_{i}^{+}|<p$, and consider an execution of the algorithm that consider vertices $v\in L_{i-1}$ in an arbitrary order.
Observe first that for any $v\in L_{i-1}$, we have $d_{M}(T(v))<2t$. Indeed, in particular when $d_{M}(T(v))\geq t$, we cannot have $X(v)=\emptyset$ as then $d_{M}(T(v))=d_{M}(v)\leq t$. We can thus take any $j\in X(v)$, and we have $d_{M}(T(v))\leq d_{M}\left(\{v\}\cup\bigcup_{k\in X(v)\setminus\{j\}}T(v_{k})\right)+d_{M}\left(T(v_{j})\right)<2t$ by minimality of $X(v)$, and as $T(v_{j})\in\mathcal{T}_{i-1}^{-}$. This directly implies that at any time $\mathcal{T}_{i-1}$ is a $t$-uniform partition of the treated vertices.
Now, let us prove the following invariant called $\Pi_{1}$. At any point when considering all the $v\in L_{i-1}$ and constructing $\mathcal{T}_{i-1}$, $\mathcal{T}_{i-1}^{+}$, and $\mathcal{T}_{i-1}^{-}$:
  • If a root $r$ of a tree in $\mathcal{T}_{i-1}^{-}$ has a parent $u$ in $G-M$, then either $r\in L_{i-1}$ and $u\in L_{i-2}$ ($u$ is not treated), or $r\in L_{i}$ and $u\in L_{i-1}$ with $u$ being either not treated yet, or the root of a tree in $\mathcal{T}_{i-1}^{+}$.
  • $\mathcal{T}_{i-1}^{-}$ is a set of disjoint non-adjacent subtrees.
For the first property, let $r$ be the root of a tree $T\in\mathcal{T}_{i-1}^{-}$ which has a parent $u$ in $G-M$. If $r\in L_{j}$ for $j>i$, by hypothesis $\mathcal{T}_{i}$ is a $t$-uniform partition of $F_{i}$, then $u$ is the root of a tree $T\in\mathcal{T}_{i}^{+}$, and $T\in\mathcal{T}_{i-1}^{+}$ as we never remove trees from $\mathcal{T}_{i}^{+}$. If $r\in L_{i-1}$, then $u\in L_{i-2}$ and we are done. Finally, assume that $r\in L_{i}$. If $u$ is not treated then we are done. Otherwise, as $u$ is treated and $r$ is still the root of a tree $T\in\mathcal{T}_{i-1}^{-}$, it implies that $X(u)\subsetneq Y^{-}(u)$, and thus that $T(u)$ was added to $\mathcal{T}_{i-1}^{+}$, satisfying the required condition.
For the second property, first note that the vertices that are in a tree of $\mathcal{T}_{i-1}^{-}$ are either vertices of $L_{i-1}$, or these vertices were in a tree of $\mathcal{T}_{i}^{-}$. By induction there is no edge between two distinct trees in $\mathcal{T}_{i}^{-}$, we are thus left with the case of an edge $uv$ with $u\in L_{i}$ and $v\in L_{i-1}$, $u$ being the child of $v$. If $u$ and $v$ are in different trees of $\mathcal{T}_{i-1}$, it is either because $u\notin Y^{-}(v)$ or because $X(v)\subset Y^{-}(v)\setminus\{u\}$. In the former case, $u$ belongs to a tree of $\mathcal{T}^{+}_{i}\subseteq\mathcal{T}^{+}_{i-1}$, and in the latter case that is $T(v)$ which belongs to $\mathcal{T}^{+}_{i-1}$.
Let us now finish the proof of 3.28. If the algorithm stops in case 2, then observe that by definition $F$ is downward-closed. Moreover, the output $\mathcal{T}_{i-1}$ is indeed a $t$-uniform partition of $F$. Indeed, Property of item 2 of 3.24 holds as, by item 1 of property $\Pi_{1}$, if a root $r$ of a tree in $\mathcal{T}_{i-1}^{-}$ has a parent $u$ in $G-M$, then $u\in L_{i-2}$ (and thus does not belong to $F$), or $u\in L_{i-1}$ and $u$ is not treated (so $u$ does not belong to $F$), or $u$ is the root of a tree in $\mathcal{T}_{i-1}^{+}$. Thus, in Case 2 we obtain $\mathcal{T}_{i-1}$, a $t$-uniform partition of $F$ with $\mathcal{T}_{i-1}^{+}=p$. By 3.26, we can find in polynomial time a $\widetilde{K_{t,t}}$ and output it.
Finally, if the algorithm does not stop, it remains to prove that $\mathcal{T}_{i-1}$ is a $t$-uniform partition of $F_{i-1}$ with $|\mathcal{T}_{i-1}^{+}|<p$. The inequality $|\mathcal{T}_{i-1}^{+}|<p$ holds as we did not fall into Case 2. Property of item 2 of 3.24 holds because we proved that for any $v\in L_{i-1}$, $d_{M}(T(v))<2t$, and we did not fall into Case 1. Property of item 3 of 3.24 holds as by item 1 of property $\Pi_{1}$, if a root $r$ of a tree in $\mathcal{T}_{i-1}^{-}$ has a parent $u$ in $G-M$, then $u\in L_{i-2}$ (and thus does not belong to $F_{i-1}$), or $u\in L_{i-1}$ and is not treated yet (which is not possible as all vertices of $L_{i-1}$ are treated in this case), or $u$ is the root of a tree in $\mathcal{T}_{i-1}^{+}$. ∎
Now that 3.28 is proved, 3.27 directly holds by induction on $i$. ∎
Remark 3.29.
Observe that in Line 5, all required conditions listed in 3.27 to use the method Partitionning-Algorithm  are fulfilled, as in particular as (KR3) does not apply, we get $d_{M}(v)\leq 2d_{r}\leq t$ for any $v\in V(G)-M$.
3.8. Final step when $G-M$ admits a $t$-uniform partition with a small number of large-degree parts
In this section, we consider Line 13 of Algorithm 1 where we found a $t$-uniform partition $\mathcal{T}$ as required. To bound $|V(G)|$ the algorithm moves some sets $Z_{1}(\mathcal{T})$ and $Z_{2}(\mathcal{T})$ (as defined below) to $M$ (to get a slightly larger set $\tilde{M}$ and tries to apply Rule (KR4) or Rule (KR5) on $(G,k,\tilde{M},\mathcal{H})$. If it not possible, we can prove that $|V(G)-M|$ is small.
Definition 3.30.
Given a $t$-uniform partition $\mathcal{T}$ of $G-M$, we define the set $Z_{1}(\mathcal{T})$ as the roots of the trees in $\mathcal{T}^{+}$, and $Z_{2}(\mathcal{T})$ as the set of vertices $v$ of $G-M$ having three edge disjoint paths $P_{1},P_{2},P_{3}$ in $G-M$ linking them to vertices of $Z_{1}(\mathcal{T})$.
Lemma 3.31.
@@ 3.8. Final step when $G-M$ admits a $t$-uniform partition with a small number of large-degree parts
Given an instance $(G,k,M,\mathcal{H})$, an integer $t$, and $\mathcal{T}$ a $t$-uniform partition of $G-M$, we have $|Z_{1}(\mathcal{T})|+|Z_{2}(\mathcal{T})|\leq 2|\mathcal{T}^{+}|$, and by denoting $\tilde{M}=M\cup Z_{1}(\mathcal{T})\cup Z_{2}(\mathcal{T})$, we have the following properties:
  • Rule (KR1) and Rule (KR2) do not apply on $(G,k,\tilde{M},\mathcal{H})$.
@@
We denote $Z_{1}=Z_{1}(\mathcal{T})$ and $Z_{2}=Z_{2}(\mathcal{T})$. The bound $|Z_{1}|\leq|\mathcal{T}^{+}|$ is trivial. To bound $|Z_{2}|$, let us define the following graph $G_{Z}$. The graph $G_{Z}$ is obtained from $G-M$ by iteratively deleting leaves $\ell$ if $\ell\notin Z_{1}\cup Z_{2}$, or contracting every edge $zx$ with $z\in Z_{1}\cup Z_{2}$ and $x\notin Z_{1}\cup Z_{2}$. Clearly $G_{Z}$ is a tree with vertex set $Z_{1}\cup Z_{2}$, and where $d_{G_{Z}}(v)\geq 3$ for every $v\in Z_{2}$. Indeed, the three edge disjoint paths leaving $v$ do not completely vanish while constructing $G_{Z}$. As the number of vertices of degree at least three in a tree is bounded by the number of leaves, we get $|Z_{2}|\leq|Z_{1}|\leq|\mathcal{T}^{+}|$.
The non applicability of Rule (KR1) and Rule (KR2) on $(G,k,\tilde{M},\mathcal{H})$ immediately follows from the non applicability of Rule (KR1) and Rule (KR2) on $(G,k,M,\mathcal{H})$.
It then remains to bound $d_{\tilde{M}}(T)$ for $T$ a connected component of $G-\tilde{M}$. Let $T$ be a connected component of $G-\tilde{M}$. We have $T\subseteq T^{\prime}$ for some $T^{\prime}\in\mathcal{T}$, as each edge between trees of $\mathcal{T}$ contains a vertex from $Z_{1}(\mathcal{T})\subseteq\tilde{M}$, and as $\mathcal{T}$ is a $t$-uniform partition, we get $d_{M}(T)\leq d_{M}(T^{\prime})\leq 2t$. Moreover, notice that $d_{\tilde{M}}(T)\leq d_{M}(T)+|\{vz\in E(G)\ \mid\ v\in T\text{ and }z\in Z_{1}\cup Z_{2}\}|$. Thus, assume by contradiction that $d_{\tilde{M}}(T)\geq 2t+3$, and thus there are three vertices $v_{1},v_{2},v_{3}$ in $T$ such that $N(v_{i})\cap(Z_{1}\cup Z_{2})\neq\emptyset$. Let $v$ be the vertex of $T$ such that the three edges $v_{1}z_{1}$, $v_{2}z_{2}$, and $v_{3}z_{3}$, where the vertices $z_{1},z_{2},z_{3}$ are distinct (but not necessarily the vertices $v_{1},v_{2},v_{3}$). The vertex $v$ at the intersection of the three paths among $v_{1},v_{2},v_{3}$ should be in $Z_{2}$, a contradiction. ∎
@@
Given an instance $(G,k,\tilde{M},\mathcal{H})$ such that Rules (KR1), (KR2), and (KR4) do not apply, and such that for any $T$ connected component of $G-\tilde{M}$, Rule (KR5) does not apply on $T$ and $d_{\tilde{M}}(T)\leq 2t+2$. Then, $|G-\tilde{M}|\leq p_{4}(r,t)|\tilde{M}|$ with $p_{4}(r,t)=(2t+4)f_{2}(r,2t+2,p_{1}(r,2t+2))$.
@@
Let $\mathcal{MC}$ be the set of connected components of $G-\tilde{M}$. Let $T\in\mathcal{MC}$. As Rule (KR1), (KR2), and (KR4) do not apply, and Rule (KR5) does not apply on $T$, by 3.20 we get $|T|\leq p_{1}(r,\operatorname{\mu}(T))\leq p_{1}(r,2t+2)$. Now, as Rule (KR4) does not apply in particular with $\mathcal{MC}$, by the second result of 3.13 with $A=\tilde{M}$, $x=2t+4$, $p=2t+2$ and $m=p_{1}(r,2t+2)$, we get that if $|\mathcal{MC}|>xf_{2}(r,p,m)|\tilde{M}|$, then we would have $x$ connected components of $\mathcal{MC}$ having the same neighborhood $X$ in $\tilde{M}$. Then we would have $X\neq\emptyset$ as otherwise Rule (KR1) would applied, and as $|X|\leq p=2t+2$ and $x\geq|X|+2$, (KR4) could be applied, a contradiction with the hypothesis. Thus, we get $|G-\tilde{M}|\leq(2t+4)f_{2}(r,2t+2,p_{1}(r,2t+2))|\tilde{M}|$. ∎
@@
Let $(G,k,M,\mathcal{H})$ be an input of $(r,\mathcal{G})$-Ann-FVS. If $A(G,k,M,\mathcal{H})$ reaches Line 23, then $|G-M|=\mathcal{O}(p_{3}(r,t)p_{4}(r,t)|M|)$.

### v2

@@ 3.7. Attempting to build a $t$-uniform partition of GMG-M
Regarding the running time, note that the set $\mathcal{T}^{\prime}$ can be computed in polynomial time, and finding the subset $X$ and $\mathcal{T}^{\prime}_{X}$ whose existence is proved by 3.11 can again be done in polynomial time by listing the neighborhoods in $M$ of the trees in $\mathcal{T}^{\prime}$. ∎
We are now finally ready to describe the aforementioned Partitionning-Algorithm  , which is done with the following lemma.
Lemma 3.25 (Partitionning lemma).
Let $t$ be an integer. There exists a polynomial time algorithm that, given an instance $(G,k,M,\mathcal{H})$ of $(r,\mathcal{G})$-Ann-FVS where none of Rules (KR1), (KR2), (KR4) applies and such that $d_{M}(v)<t$ for every $v\in V(G-M)$, returns one of the following:
  1. Case 1
    a subtree $T$ of the forest GMG-M with $|T|>p_{1}(r,\operatorname{\mu}(T))$
  2. Case 2
    a $\widetilde{K_{t,t}}$, or
  3. Case 3
    a $t$-uniform partition $\mathcal{T}$ of the forest GMG-M with $|\mathcal{T}^{+}|<p_{3}(r,s,t)|M|$, where $p_{3}$ is defined in 3.24.
Observe that in Case 1 and Case 2 above, the rules (KR5) and (BR2) respectively apply, which will allow us to make progress even if we cannot find a partition.
Proof.
In this proof we set $p=p_{3}(r,s,t)|M|$. Let $h$ be the maximum depth of a leaf in GMG-M and let $L_{0}$ be the set of roots of the connected components of GMG-M. For every $i\in\left\{1,\dots,h\right\}$, we define $L_{i}$ as the set of vertices of depth $i$ in the trees of the forest GMG-M and $F_{i}$ as the forest induced by the set of vertices with depth at least $i$, i.e., $F_{i}=G\left[\bigcup_{j=i}^{h}L_{j}\right]$. Observe that $F_{0}=G-M$.
The algorithm we describe here considers vertices by decreasing depth. First it construct a (trivial) partition of the forest $F_{h}$ induced by the vertices of maximum depth. Observe that two vertices of the same depth cannot share an edge, so $F_{h}$ an edgeless graph. In order to construct the partition we then have no choice but to take each connected component (containing only one vertex) as a tree in $\mathcal{T}$, and more precisely in $\mathcal{T}_{h}^{-}$ as for $v\in V(F_{h})$ we have $d_{M}(v)<t$ by hypothesis, and so $\mathcal{T}_{h}^{+}=\emptyset$. The collection of trees $\mathcal{T}_{h}$ is then a $t$-uniform partition as there is no edge between two trees of $\mathcal{T}_{h}$, and each tree have degree in $M$ at most $2t$ as wanted (even $t$, as noted above).
The second stage of the algorithm consists in repeatedly applying a routine to try to extend a $t$-uniform partition of the vertices at depth more than $i$ to a $t$-uniform partition of the vertices at depth more than i1i-1, for decreasing $i$. Possibly this extension will fail, in which case we show that we can return a large subtree of GMG-M or a $\widetilde{K_{t,t}}$. If the successive extensions succeed up to partitioning V(GM)V(G-M), we show that the result is a $t$-uniform partition. The routine is the following.
  1. (1)
    Input: a $t$-uniform partition $\mathcal{T}_{i}$ of $F_{i}$ such that $|\mathcal{T}_{i}^{+}|<p$.
  2. (2)
    Output: a $t$-uniform partition $\mathcal{T}_{i-1}$ of $F_{i-1}$ with $|\mathcal{T}_{i-1}^{+}|<p$, or a $\widetilde{K_{t,t}}$, or a subtree of GMG-M with more than $p_{1}(r,\operatorname{\mu}(T))$ vertices. The fact that the routine described here indeed corresponds to this specification is given in 3.26.
  3. (3)
    Initialization: we start with $\mathcal{T}_{i-1}^{+}=\mathcal{T}_{i}^{+}$ and $\mathcal{T}_{i-1}^{-}=\mathcal{T}_{i}^{-}$. The collection $\mathcal{T}_{i-1}$ is always defined as the union $\mathcal{T}_{i-1}^{-}\cup\mathcal{T}_{i-1}^{+}$.
  4. (4)
    For every $v\in L_{i-1}$:
    1. (LABEL:enumiie:for.1)
      Let $v_{1},\dots,v_{l}$ denote the children of $v$, which all belong to $L_{i}$. Observe that by definition of a $t$-uniform partition of $F_{i}$, each such vertex $v_{j}$ is the root of some tree $T(v_{j})$ in $\mathcal{T}_{i}$. Informally, the goal will be to try to group some of the trees of the form $T(v_{j})\in\mathcal{T}_{i}^{-}$ to construct a tree to add in $\mathcal{T}_{i-1}^{+}$.
    2. (LABEL:enumiie:for.2)
      Let $Y^{-}(v)=\{j\mid T(v_{j})\in\mathcal{T}_{i}^{-}\}$.
    3. (LABEL:enumiie:for.3)
      If $d_{M}({v}\cup\bigcup_{j\in Y^{-}(v)}T(v_{j}))<t$, set $X(v)=Y^{-}(v)$. Otherwise, define $X(v)$ as an inclusion-wise minimal subset of $Y^{-}(v)$ such that $d_{M}\left(\{v\}\cup\bigcup_{j\in X(v)}T(v_{j})\right)\geq t$.
    4. (LABEL:enumiie:for.4)
      Let $T(v)=\{v\}\cup\bigcup_{j\in X(v)}T(v_{j})$.
    5. (LABEL:enumiie:for.5)
      If $|T(v)|>p_{1}(r,\operatorname{\mu}(T))$ output the tree $T(v)$ and stop.
    6. (LABEL:enumiie:for.6)
      If $d_{M}(T(v))<t$, remove from $\mathcal{T}_{i-1}^{-}$ all the $T(v_{j})$ for $j\in X(v)$, and add $T(v)$ to $\mathcal{T}_{i-1}^{-}$. Vertex $v$ has been dealt with: continue in line (4) to the next choice of a vertex in $L_{i-1}$.
    7. (LABEL:enumiie:for.7)
      Otherwise ($d_{M}(T(v))\geq t$), remove from $\mathcal{T}_{i-1}^{-}$ all the $T(v_{j})$ for $j\in X(v)$, and add $T(v)$ to $\mathcal{T}_{i-1}^{+}$.
    8. (LABEL:enumiie:for.8)
      If $|\mathcal{T}_{i-1}^{+}|<p$, continue in line (4) to the next choice of vertex of $L_{i-1}$.
    9. (LABEL:enumiie:for.9)
      Otherwise, $|\mathcal{T}_{i-1}^{+}|=p$. In this case we apply 3.24 to the instance $(G,k,M,\mathcal{H})$ with partition $\mathcal{T}_{i-1}$ of the forest $F=\bigcup_{T\in\mathcal{T}_{i-1}}T$. As we will show, $\mathcal{T}_{i-1}$ is a $t$-uniform partition of $F$ so this call results in a $\widetilde{K_{t,t}}$. Output the obtained $\widetilde{K_{t,t}}$ and stop.
  5. (5)
    Return $\mathcal{T}_{i-1}$.
The correctness of the above routine is given by the following claim.
Claim 3.26.
The routine above is correct, i.e., given an input as described in line (1) above, the routine returns an output as specified in line (2). Moreover, the computation can be done in polynomial time.
Proof.
Let us consider as input $\mathcal{T}_{i}$ a $t$-uniform partition of $F_{i}$ with $i\geq 1$ such that $|\mathcal{T}_{i}^{+}|<p$. We consider an execution of the routine on an input as described in line (1) and that considers vertices $v\in L_{i-1}$ in an arbitrary order.
There are 3 places where the routine returns a value. If this happens at line (4)(e:for.5), the returned value is clearly a subtree of GMG-M with more than $p_{1}(r,\operatorname{\mu}(T))$ vertices. So in the following we may assume that we are not in this case and so for every vertex $v\in L_{i-1}$ considered so far, $|T(v)|\leq p_{1}(r,\operatorname{\mu}(T))$.
If we return at line (4)(e:for.9) then the output is a $\widetilde{K_{t,t}}$ according to 3.24, supposing we gave an appropriate input to the algorithm of that lemma. By construction $F$ is downwards-closed. To show that $\mathcal{T}_{i-1}$ is a $t$-uniform partition of $F$, let us consider the 3 items in 3.22. Item 1 holds by construction. For every tree of $\mathcal{T}_{i-1}$ of the form $T(v)$, either $v\in V(F_{i})$ and so the bounds of Item 2 hold since $\mathcal{T}_{i}$ is a uniform $t$-partition, or $v\in L_{i-1}$ and as noted above we assume $|T(v)|\leq p_{1}(r,\operatorname{\mu}(T))$. It remains to show that for such a vertex $v$ we have $d_{M}(T(v))\leq 2t$. To do so, notice that when $d_{M}(T(v))\geq t$, then $X(v)\neq\emptyset$ as otherwise we would get $d_{M}(T(v))=d_{M}(v)<t$, a contradiction. We can thus consider any $j\in X(v)$, and observe that $d_{M}(T(v))\leq d_{M}\left(\{v\}\cup\bigcup_{k\in X(v)\setminus\{j\}}T(v_{k})\right)+d_{M}\left(T(v_{j})\right)<2t$ by minimality of $X(v)$, and as $T(v_{j})\in\mathcal{T}_{i-1}^{-}$. To show Item 3, we will use the following invariant.
  • At any time during the execution of the routine, if a root $r$ of a tree in $\mathcal{T}_{i-1}^{-}$ has a parent $u$ in GMG-M, then either $r\in L_{i-1}$ and $u\in L_{i-2}$, or $r\in L_{i}$, $u\in L_{i-1}$, and $u$ either has not been considered so far or is the root of a tree in $\mathcal{T}_{i-1}^{+}$.
The proof of the invariant is the following. Let $r$ be the root of a tree $T\in\mathcal{T}_{i-1}^{-}$ which has a parent $u$ in GMG-M. If $r\in L_{j}$ for $j>i$, by hypothesis $\mathcal{T}_{i}$ is a $t$-uniform partition of $F_{i}$, then $u$ is the root of a tree $T\in\mathcal{T}_{i}^{+}$, and $T\in\mathcal{T}_{i-1}^{+}$ as we never remove trees from $\mathcal{T}_{i}^{+}$. If $r\in L_{i-1}$, then $u\in L_{i-2}$ and we are done. Finally, assume that $r\in L_{i}$. If $u$ has not been considered so far then we are done. Otherwise, as $u$ has been considered and $r$ is still the root of a tree $T\in\mathcal{T}_{i-1}^{-}$, it implies that $X(u)\neq Y^{-}(u)$, and thus that $T(u)$ was added to $\mathcal{T}_{i-1}^{+}$, satisfying the required condition. This concludes the proof of the invariant.
We can now show that Item 3 of 3.22 holds for $\mathcal{T}_{i-1}$. By the first item of the invariant, if a root $r$ of a tree in $\mathcal{T}_{i-1}^{-}$ has a parent $u$ in GMG-M, then $u\in L_{i-2}$ (and thus does not belong to $F$), or $u\in L_{i-1}$ and $u$ has not been considered so far (so $u$ does not belong to $F$), or $u$ is the root of a tree in $\mathcal{T}_{i-1}^{+}$. This shows that $\mathcal{T}_{i-1}$ is indeed a $t$-uniform partition of $F$ hence the call to the algorithm of 3.24 is valid. This concludes the proof that in line (4)(e:for.9) the routine returns a $\widetilde{K_{t,t}}$.
[…]
Now that 3.26 is proved, 3.25 directly holds by induction on $i$, using $\mathcal{T}_{h}$ for the base case. ∎
Remark 3.27.
Observe that we call Partitionning-Algorithm  of 3.25 at line 5 of Algorithm 1, all required conditions on the input are fulfilled, as in particular as (KR3) does not apply, we get $d_{M}(v)\leq 2d(r)\leq t$ for any $v\in V(G-M)$.
3.8. Final step when GMG-M admits a $t$-uniform partition with a small number of large-degree parts
[…]
Definition 3.28.
Given a $t$-uniform partition $\mathcal{T}$ of GMG-M, we define the set $Z_{1}(\mathcal{T})$ as the roots of the trees in $\mathcal{T}^{+}$, and $Z_{2}(\mathcal{T})$ as the set of vertices $v$ of GMG-M having three edge-disjoint paths $P_{1},P_{2},P_{3}$ in GMG-M linking them to vertices of $Z_{1}(\mathcal{T})$.
@@ 3.8. Final step when GMG-M admits a $t$-uniform partition with a small number of large-degree parts
Given an instance $(G,k,M,\mathcal{H})$, an integer $t$, and $\mathcal{T}$ a $t$-uniform partition of GMG-M, we have $|Z_{1}(\mathcal{T})|+|Z_{2}(\mathcal{T})|\leq 2|\mathcal{T}^{+}|$, and by denoting $\tilde{M}=M\cup Z_{1}(\mathcal{T})\cup Z_{2}(\mathcal{T})$, we have the following properties:
[…]
@@
We denote $Z_{1}=Z_{1}(\mathcal{T})$ and $Z_{2}=Z_{2}(\mathcal{T})$. The bound $|Z_{1}|\leq|\mathcal{T}^{+}|$ is trivial. To bound $|Z_{2}|$, let us define the following minor $G_{Z}$ of GMG-M. The graph $G_{Z}$ is obtained from GMG-M by iteratively deleting leaves $\ell$ if $\ell\notin Z_{1}\cup Z_{2}$, or contracting every edge $zx$ with $z\in Z_{1}\cup Z_{2}$ and $x\notin Z_{1}\cup Z_{2}$. Clearly $G_{Z}$ is a forest with vertex set $Z_{1}\cup Z_{2}$. Furthermore, for every $v\in Z_{2}$, the three edge-disjoint paths linking $v$ to $Z_{1}$ are actually vertex-disjoint and end up on different vertices of $Z_{1}$, hence $d_{G_{Z}}(v)\geq 3$. As the number of vertices of degree at least three in a tree is bounded by the number of leaves, we get $|Z_{2}|\leq|Z_{1}|\leq|\mathcal{T}^{+}|$.
[…]
It then remains to bound $d_{\tilde{M}}(T)$ for $T$ a connected component of $G-\tilde{M}$. Let $T$ be a connected component of $G-\tilde{M}$. We have $T\subseteq T^{\prime}$ for some $T^{\prime}\in\mathcal{T}$, as each edge between trees of $\mathcal{T}$ contains a vertex from $Z_{1}(\mathcal{T})\subseteq\tilde{M}$, and as $\mathcal{T}$ is a $t$-uniform partition, we get $d_{M}(T)\leq d_{M}(T^{\prime})\leq 2t$. Moreover, notice that $d_{\tilde{M}}(T)\leq d_{M}(T)+|\{vz\in E(G)\ \mid\ v\in T\text{ and }z\in Z_{1}\cup Z_{2}\}|$. Thus, assume by contradiction that $d_{\tilde{M}}(T)\geq 2t+3$, and thus there are three (non-necessarily distinct) vertices $v_{1},v_{2},v_{3}$ in $T$ and three distinct vertices $z_{1},z_{2},z_{3}$ in $Z_{1}\cup Z_{2}$ such that $v_{1}z_{1}$, $v_{2}z_{2}$, and $v_{3}z_{3}$ are edges. Then $T$ contains a vertex $v$ linked by 3 edge-disjoint paths to $Z_{1}\cup Z_{2}$. As such a path ending in $Z_{2}$ could be prolonged towards $Z_{1}$ while staying edge disjoint from the others, such a vertex $v$ should belong to $Z_{2}$, a contradiction. ∎
@@
[…]
@@
[…]
@@
Let $(G,k,M,\mathcal{H})$ be an instance of $(r,\mathcal{G})$-Ann-FVS. If on this input Algorithm 1 reaches Line 23, then $|G-M|=\mathcal{O}(p_{3}(r,t)p_{4}(r,t)|M|)$.